TIME RESOLVED INFRARED DIODE LASER SPECTROSCOPY OF THE ν_{1} BAND OF THE CoCO RAD-ICAL PRODUCED BY THE ULTRAVIOLET PHOTOLYSIS OF Co(CO)_{3}NO.

\underline{TOSHIHIDE HIKIDA}, SEIKI IKEDA, KEIICHI TANAKA and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

Rovibrational transitions of the ν_{1} band (C-O stretch) of the CoCO radical produced by the ultraviolet photolysis of Co(CO)_{3}NO were observed in 1949-1991 cm^{−1} region by the time resolved infrared diode laser spectroscopy. The electronic ground state was confirmed to be X^{2}Δ_{i} by the analysis as predicted by the density functional theory calculation. ^{a} More than 140 absorption lines were assigned to the Ω = 5/2 spin component of the ν_{1} fundamental band in the X^{2}Δ_{i} state. No lines of the Ω = 3/2 component were observed due to the large spin-orbit interaction constant A (−500 cm^{−1} ). Effective molecular constants for the Ω = 5/2 spin component, including the band origin ν_{0} and the rotational constant B, were derived from observed transitions. Hot band lines originated from the ν_{2} (Co-C-O bending) and ν_{3} (Co-C stretch) vibrational excited states were observed. Absorption lines for the ν_{2} hot band were split into two components because of the vibronic interaction between the ν_{2} state and other nearby electronic excited states. The effective equilibrium rotational constant B_{e} and the vibration rotation constants, α_{1}, α_{2} and α_{3} were determined from the analysis of the fundamental band and hot bands.

aM. Zhou, L. Andrews, J.P hys.Chem. A 102, 10250 (1998)